# When Should a Model Change Its Mind? An Energy-Based Theory and Regularizer for Concept Drift in Electrocardiogram (ECG) Signals

### v1

@@ V Results
This section evaluates the behavior of unimodal and multimodal ECG models under physiologic perturbations and quantifies the effect of Energy-Constrained Representation Learning (ECRL). All clean and perturbed accuracies, as well as robustness degradation metrics, are summarized in Tables I, II, and III.
@@ V-A Unimodal Baselines
Table I shows clear differences in sensitivity among the three unimodal models. The temporal 1D-CNN (M1) experiences the most severe robustness degradation: perturbed accuracy drops from $0.9355$ to $0.3925$ (robustness degradation =0.5430=0.5430). This confirms that temporal morphology is highly sensitive to benign physiologic variations such as mild amplitude or rate shifts, consistent with PECT’s prediction that raw temporal encoders exhibit large latent movement when physiologic energy changes.
@@ V-B Effect of BIT+PECT on Unimodal Models
TABLE I: Unimodal ECG Models: Baseline vs. BIT+PECT.
@@ V-C Multimodal Fusion Baselines
Table II highlights that multimodal fusion, while beneficial for clean accuracy, magnifies instability under physiologic perturbation. The strongest clean-performing model (M7: 1D+2D+TR) achieves a clean accuracy of $0.9600$, yet its perturbed accuracy drops to $0.7258$ (robustness degradation =0.2342=0.2342). Similar degradation is observed in the other multimodal architectures: M5 falls to $0.2527$ perturbed accuracy, and M6 drops from $0.9400$ to $0.7258$.
@@
TABLE II: Multimodal Late-Fusion ECG Models: Baseline vs. BIT+PECT.
@@ V-E Percentage Improvements
Table III quantifies relative changes across all seven models. The largest gains occur in perturbed accuracy, notably:
@@
TABLE III: Percentage improvement introduced by BIT+PECT.
@@ VI-A When Should a Model Change Its Mind?
Within the framework of Physiologic Energy Conservation Theory (PECT), the title question can now be stated precisely. A model should change its prediction only when representation drift is both decision-relevant and physiologically justified. PECT formalizes this requirement through an energy–drift proportionality principle that links latent geometry directly to interpretable physical variation in the underlying signal.
Two regimes emerge. When the physiologic energy shift $|\Delta E_{\mathrm{phys}}|$ is small but the fused latent displacement $\|\delta z_{\mathrm{fused}}\|$ is large, the model should not change its prediction. This corresponds to virtual drift: benign, label-preserving variation that induces spurious movement in latent space. In contrast, when latent displacement exceeds the classifier margin and scales proportionally with a substantial energy change, the model should change its prediction, reflecting true concept change.
@@
$\|\delta z_{\mathrm{fused}}\|\geq\gamma(x)\quad\text{and}\quad\|\delta z_{\mathrm{fused}}\|\propto|\Delta E_{\mathrm{phys}}|.$
Thus, decision boundaries should be crossed only when representation drift is both margin-violating and energetically justified. This provides a principled, biophysically grounded criterion for distinguishing true concept change from benign variability in dynamic physiologic signals.
@@ VI-B Interpretation of Experimental Results
The experimental results demonstrate that physiologic energy provides a meaningful and interpretable lens for understanding representation instability in multimodal ECG models and, more broadly, for reasoning about concept drift in dynamic signals. Unlike adversarial or noise-based perturbations, physiologic variations arise from predictable electromechanical processes in the heart and correspond to label-preserving virtual drift. PECT formalizes this structure by asserting that latent drift should scale proportionally with changes in physiologic energy. When this proportionality is violated, even mild variations can induce disproportionate movement in latent space, destabilizing multimodal fusion and manifesting as spurious concept-drift events.
@@ VII Conclusion
This work introduced Physiologic Energy Conservation Theory (PECT), an energy-based framework for interpreting concept drift in multimodal ECG deep learning. PECT formalizes the principle that latent representation drift should scale proportionally with physiologic energy variation and explains why multimodal fusion—despite improving clean accuracy—can destabilize predictions under benign perturbations.
Building on this theory, we proposed Energy-Constrained Representation Learning (ECRL), a lightweight training framework that enforces energy-proportional latent geometry without modifying model architecture or increasing inference cost. Experiments across seven unimodal and multimodal models demonstrate that ECRL substantially reduces drift, stabilizes multimodal fusion, and improves robustness while preserving competitive clean performance.

### v2

@@ V Results
This section evaluates the behavior of unimodal and multimodal ECG models under physiologic perturbations and quantifies the effect of Energy-Constrained Representation Learning (ECRL). All clean and perturbed accuracies, as well as robustness degradation metrics, are summarized in Tables II, III, and IV.
@@ V-A Unimodal Baselines
Table II shows clear differences in sensitivity among the three unimodal models. The temporal 1D-CNN (M1) experiences the most severe robustness degradation: perturbed accuracy drops from $0.9355$ to $0.3925$ (robustness degradation =0.5430=0.5430). This confirms that temporal morphology is highly sensitive to benign physiologic variations such as mild amplitude or rate shifts, consistent with PECT’s prediction that raw temporal encoders exhibit large latent movement when physiologic energy changes.
@@ V-B Effect of BIT+PECT on Unimodal Models
TABLE II: Unimodal ECG Models: Baseline vs. BIT+PECT.
@@ V-C Multimodal Fusion Baselines
Table III highlights that multimodal fusion, while beneficial for clean accuracy, magnifies instability under physiologic perturbation. The strongest clean-performing model (M7: 1D+2D+TR) achieves a clean accuracy of $0.9600$, yet its perturbed accuracy drops to $0.7258$ (robustness degradation =0.2342=0.2342). Similar degradation is observed in the other multimodal architectures: M5 falls to $0.2527$ perturbed accuracy, and M6 drops from $0.9400$ to $0.7258$.
@@
TABLE III: Multimodal Late-Fusion ECG Models: Baseline vs. BIT+PECT.
@@ V-E Percentage Improvements
Table IV quantifies relative changes across all seven models. The largest gains occur in perturbed accuracy, notably:
@@
TABLE IV: Percentage improvement introduced by BIT+PECT.
@@ VI-A When Should a Model Change Its Mind?
Within the framework of Physiologic Energy Conservation Theory (PECT), the title question can now be stated precisely. A model should change its prediction only when representation drift is both decision-relevant and physiologically justified. PECT formalizes this requirement through an energy–drift consistency principle that links latent geometry directly to interpretable physical variation in the underlying signal.
Two regimes emerge. When the physiologic energy shift $|\Delta E_{\mathrm{phys}}|$ is small but the fused latent displacement $\|\delta z_{\mathrm{fused}}\|$ is large, the model should not change its prediction. This corresponds to virtual drift: benign, label-preserving variation that induces spurious movement in latent space. In contrast, when latent displacement exceeds the classifier margin and remains consistent with a substantial energy change, the model should change its prediction, reflecting true concept change.
@@
$\|\delta z_{\mathrm{fused}}\|\geq\gamma(x)\quad\text{and}\quad\|\delta z_{\mathrm{fused}}\|\ \text{is consistent with}\ |\Delta E_{\mathrm{phys}}|.$
Thus, decision boundaries should be crossed only when representation drift is both margin-violating and energetically consistent. This provides a principled, biophysically grounded criterion for distinguishing true concept change from benign variability in dynamic physiologic signals.
@@ VI-B Interpretation of Experimental Results
The experimental results demonstrate that physiologic energy provides a meaningful and interpretable lens for understanding representation instability in multimodal ECG models and, more broadly, for reasoning about concept drift in dynamic signals. Unlike adversarial or noise-based perturbations, physiologic variations arise from predictable electromechanical processes in the heart and correspond to label-preserving virtual drift. PECT formalizes this structure by asserting that latent drift should be energy-consistent with changes in physiologic energy. When this relationship is violated, even mild variations can induce disproportionate movement in latent space, destabilizing multimodal fusion and manifesting as spurious concept-drift events.
@@ VII Conclusion
This work introduced Physiologic Energy Conservation Theory (PECT), an energy-based framework for interpreting concept drift in multimodal ECG deep learning. PECT formalizes the principle that latent representation drift should remain consistent with physiologic energy variation and explains why multimodal fusion (despite improving clean accuracy) can destabilize predictions under benign perturbations.
Building on this theory, we proposed Energy-Constrained Representation Learning (ECRL), a lightweight training framework that enforces energy-consistent latent geometry without modifying model architecture or increasing inference cost. Experiments across seven unimodal and multimodal models demonstrate that ECRL substantially reduces drift, stabilizes multimodal fusion, and improves robustness while preserving competitive clean performance.